OpenStudy (asad786):

Translate the sentence into an equation. Six less than the quotient of a number and 5 is 10. A. B. C. 5n – 6 =10 D. (n + 5) – 6 = 10

OpenStudy (asad786):

B. n/5 - 6 = 10

OpenStudy (asad786):

a. 6 - n/5 =10

OpenStudy (asad786):

@StudyGurl14

OpenStudy (luffingsails):

Hey, You have to think about this one in order. "Six less than a number Y" = Y - 6 "A quotient of a number X of 5" = X / 5 Putting it all together N/5 -6 = 10.